rayneeee:

x + y = −3 y = 2x + 2 how do I solve this equation?

ScamTheMan:

Well, this is a simple system of equations equation. Our first step is to find one value by subtracting the other one. x + y = -3 -2x + y = 2 We can multiply the top side by 2 so we can subtract the x value and solve for the y value. 2x + 2y = -6 -2x + y = 2 Solve it through by combing the top and bottom. 3y = -4 y = -4/3 We then take -4/3 and substitute it for y in the first or second equation. I'll give you the equation so you can solve for yourself: x + -4/3 = -3

rayneeee:

alr thx

ScamTheMan:

Of course <3

AZ:

Or you could have just substituted the second equation into the first haha but same stuff